another word problem The sum of two consecutive integers is 145. Find the two integers.
Answer by checkley71(8403)

X+X+1=145
2X=145-1
2X=144
X=144/2
X=72 ANSWER FOR THE FIRST INTERGER.
72+1=73 FOR THE SECOND INTERGER.
PROOF
72+73=145
145=145